# MASt3R-SLAM: Real-Time Dense SLAM with 3D Reconstruction Priors

### v1

@@ 1 Introduction
Visual simultaneous localisation and mapping (SLAM) is a foundational building block for today’s robotics and augmented reality products. With careful design of an integrated hardware and software stack, robust and accurate visual SLAM is now possible. However, SLAM is yet to become a plug-and-play algorithm as it requires hardware expertise and calibration. Even for a minimal single camera setup with no additional sensing such as an IMU, a SLAM solution that reliably provides both accurate poses and consistent dense maps in-the-wild does not exist. Achieving such a reliable dense SLAM system would open new research avenues for spatial intelligence.
Performing dense SLAM from only 2D images requires reasoning over time-varying poses and camera models, as well as 3D scene geometry. To solve such an inverse problem of large dimensionality, a variety of priors, from handcrafted to data-driven, have been proposed. Single-view priors, such as monocular depth and normals, attempt to predict geometry from a single image, but these contain ambiguities and lack consistency across views. While multi-view priors like optical flow reduce the ambiguity, decoupling pose and geometry is challenging since pixel motion depends on both the extrinsics and the camera model. Although these underlying causes may vary across time and different observers, the 3D scene remains invariant across views. Therefore, the unifying prior required to solve for poses, camera models, and dense geometry from images is over the space of 3D geometry in a common coordinate frame.
Figure 1: Reconstruction from our dense monocular SLAM system on the Burghers sequence [55]. Using the two-view predictions from MASt3R shown on the left, our system achieves globally consistent poses and geometry in real-time even without a known camera model.
@@ 2 Related Work
All systems mentioned thus far assume known intrinsic calibration. Classical automatic intrinsic calibration is possible when there are strict assumptions on scene geometry or unchanging parameters across a set of images [12], but encounters degenerate configurations and sensitivity to noise. Given an initial estimate of intrinsics, refinement via bundle adjustment can improve accuracy online [16], but this already assumes a parametric model and sufficient initialisation of all parameters. More recently, data-driven methods predict intrinsics from one or multiple images [13, 47], but are either limited in accuracy for in-the-wild SLAM or are not flexible in the camera model definition.
@@ 3.2 Pointmap Matching
By using the nonlinear least-squares form similar to [34], we can iteratively solve for updates $\mathbf{\delta p}$ by calculating analytical Jacobians $\mathbf{J}$ and using Levenberg-Marquardt:
$\mathbf{\delta p}=-\left(\mathbf{J}^{T}\mathbf{J}+\lambda\mathbf{I}\right)^{-1}\mathbf{J}^{T}\mathbf{r},\quad\mathbf{p}\leftarrow\mathbf{p}+\mathbf{\delta p}.$ (4)
This can be done separately for each point and converges for almost all valid pixels within 10 iterations as the ray image is smooth. At the end of this process, we now have initial matches ${{\mathbf{m}}}_{i,j}$. When there is no initial estimate for the projection $\mathbf{p}$, such as when tracking against a new keyframe or when matching loop closure edges, all pixels are initialised with the identity mapping. During tracking, since we always have the matches from the previous frame, we can use this as initialisation to further speed up the convergence. To handle occlusions and outliers, we also invalidate matches that have large distances in 3D space. Our matching is massively parallel on GPU and additionally can leverage the incremental nature of SLAM.
@@ 3.3 Tracking and Pointmap Fusion
While ${{\mathbf{X}}}^{k}_{f}$ instead of ${{\mathbf{X}}}^{f}_{f}$ could also be aligned to ${{\mathbf{X}}}^{k}_{k}$, with the benefit of no explicit matching required as they are pixel aligned, we found that explicit matching with ${{\mathbf{X}}}^{f}_{f}$ had improved accuracy for larger baseline scenarios. More importantly, although the 3D point error is suitable, it is easily skewed by errors in the pointmap predictions as inconsistent predictions in depth are relatively frequent. Since we ultimately fuse predictions into a single pointmap that averages out all the predictions, error in tracking degrades the keyframe’s pointmap that will also be used in the backend.
@@
This results in a similar angular error as mentioned in Eq. 3 and shown in Fig. 2, except that we now have many known correspondences and wish to find the pose that minimises all angular errors between canonical rays and corresponding predicted rays from the current frame. Since angular errors are bounded, ray-based errors are robust against outliers [29]. We also include an error term with a small weight on the difference in distances from the camera centre. This prevents the system from becoming degenerate in pure rotation cases, but the small weight avoids biasing the pose estimate in the same way that the point error does. We efficiently solve for updates to the pose using Gauss-Newton in an iteratively reweighted least-squares (IRLS) framework. We calculate analytical Jacobians of the ray and distance errors with respect to a perturbation ${\boldsymbol{\tau}}$ of the relative pose ${{\mathbf{T}}}_{kf}$. We stack the residuals, Jacobians, and weights into matrices $\mathbf{r}$, $\mathbf{J}$, and $\mathbf{W}$, respectively. We iteratively solve the linear system and update the pose via:
@@
The pointmap initially has larger errors and less confidence due to only having small baseline frames, but filtering is able to merge information from many viewpoints. We experimented with different ways of updating the canonical pointmap, and found that weighted average was best for maintaining coherence while filtering out noise. Compared to the canonical pointmap in MASt3R-SfM [10], we compute this incrementally and require transformation of the points since an additional network prediction of ${{\mathbf{X}}}^{k}_{k}$ would slow down tracking. Filtering has a rich history in SLAM, and yields the benefit of leveraging information from all frames without having to explicitly optimise for all camera poses and store all predicted pointmaps from the decoder in the backend.
@@ 3.4 Graph Construction and Loop Closure
When tracking, a new keyframe $\mathcal{K}_{i}$ is added if the number of valid matches or the number of unique keyframe pixels in ${{\mathbf{m}}}_{f,k}$ falls below a threshold $\omega_{k}$. After adding $\mathcal{K}_{i}$, a bidirectional edge to the previous keyframe $\mathcal{K}_{i-1}$is added to the edge-list $\mathcal{E}$. This constrains the estimated poses sequentially in time; however, drift can still occur. To close both small and large loops, we adapt the Aggregated Selective Match Kernel (ASMK) [45, 46] framework used by MASt3R-SfM [10] for image retrieval from encoded features. While this was previously used in a batch setting where all images are available from the start, we modify it to work incrementally. We query the database with the encoded features of $\mathcal{K}_{i}$ to obtain the top-$K$ images. Since the codebook only has tens of thousands of centroids, we found that a dense L2 distance calculation was sufficiently fast to quantise the features. If the retrieval scores are above a threshold $\omega_{r}$, we give these pairs to the MASt3R decoder and add bidirectional edges if the number of matches determined from Sec. 3.2 is above a threshold $\omega_{l}$. Lastly, we update the retrieval database by adding the new keyframe’s encoded features to the inverted file index.
@@ 3.5 Backend Optimisation
where ${{\mathbf{T}}}_{ij}={{\mathbf{T}}}_{WC_{i}}^{-1}{{\mathbf{T}}}_{WC_{j}}$. Given N keyframes, Eq. 10 forms and accumulates $14\times 14$ blocks into the $7N\times 7N$ Hessian. We solve this problem again using Gauss-Newton as in Eq. 8 but with sparse Cholesky decomposition as the system is not dense. Construction of the Hessian is made efficient through the use of analytical Jacobians and parallel reductions all implemented in CUDA. Again, a small error term on consistency in distances is added to avoid degeneracy in the pure-rotation case. At most 10 iterations of Gauss-Newton are performed for every new keyframe and optimisation terminates early upon convergence. The use of second-order information greatly speeds up the global optimisation over the alternatives, and our efficient implementation ensures that it is not the bottleneck in the overall system.
@@ 3.6 Relocalisation
If the system loses tracking due to an insufficient number of matches, relocalisation mode is triggered. For a new frame, the retrieval database is queried with a stricter threshold on the score. Once the retrieved images have a sufficient number of matches with the current frame, it is then added as a new keyframe into the graph and tracking resumes.
@@ 4.1 Camera Pose Estimation
TUM RGB-D: On the TUM dataset, we demonstrate state-of-the-art trajectory error when leveraging calibration information as shown in Tab. 1. Many of the previously best performing algorithms, such as DROID-SLAM, DPV-SLAM, and GO-SLAM, build on the foundational matching and end-to-end system proposed by DROID-SLAM. In contrast, we propose a unique system that takes an off-the-shelf two-view geometric prior and show that it can outperform all other systems while operating in real-time. Furthermore, our uncalibrated system significantly outperforms a baseline, which we denote DROID-SLAM*, that calibrates the intrinsics using GeoCalib [47] on the first image of a sequence, which is then used by DROID-SLAM. We achieve this without assuming a fixed camera model across the entire sequence, and demonstrate the value of 3D priors for dense uncalibrated SLAM over priors that solve subproblems. Our uncalibrated SLAM results are also comparable to results from other recent learned techniques such as DPV-SLAM with known calibration.
@@ 4.2 Dense Geometry Evaluation
t 7-scenes ATE Accuracy Completion Chamfer DROID-SLAM 0.049 0.141 0.048 0.094 Spann3R @20 N/A 0.069 0.047 0.058 Spann3R @2 N/A 0.124 0.043 0.084 Ours 0.047 0.089 0.085 0.087 Ours* 0.066 0.068 0.045 0.056 EuRoC ATE Accuracy Completion Chamfer DROID-SLAM 0.022 0.173 0.061 0.117 Ours 0.041 0.099 0.071 0.085 Ours* 0.164 0.108 0.072 0.090
@@ 4.3 Qualitative Results
4.4 Ablation Studies
In Tab. 5, the ray error formulation for uncalibrated tracking and backend optimisation significantly improves performance over using the 3D point error and mitigates effects of inaccurate pointmap predictions. In Tab. 5, we test different methods for updating the canonical pointmap and report the average ATE across TUM, 7-Scenes, and EuRoC. Selecting the most recent or the latest point result in lack of sufficient baseline to resolve geometry or incur drift, respectively. Given calibration, weighted fusion performs on par with selecting the full pointmap with the highest median confidence, but it achieves the lowest ATE without calibration and improves the ATE on EuRoC by 1.3cm, indicating that fusing over camera models is significant.
We compare matching techniques in Tab. 6. In general, our parallel projective matching with feature refinement achieves the best accuracy with significantly faster runtime. Performing MASt3R matching over all pixels takes around 2 seconds, while our matching takes 2ms and makes the entire system FPS nearly 40x faster. Please refer to the supplementary for a full runtime analysis of the system.
Table 4: ATE (m) for point or ray error ablation.
TUM 7-Scenes EuRoC avg
Point 0.092 0.084 0.290 0.155
Ray 0.060 0.066 0.164 0.097
Table 5: ATE (m) pointmap fusion ablation.
@@ 4.4 Ablation Studies
Median conf 0.102 0.039
Weighted fusion 0.097 0.039
Table 6: Comparison of accuracy and runtime for different matching methods.
ATE ATE Matching System
w/ calib w/o calib time (ms) FPS
$k$-d tree 0.061 0.115 40 8.8
MASt3R 0.042 0.098 2000 0.4
Ours w/o features 0.062 0.092 0.5 15.1
Ours w/ features 0.039 0.097 2 14.9

### v2

@@ 1 Introduction
Visual simultaneous localisation and mapping (SLAM) is a foundational building block for today’s robotics and augmented reality products. With careful design of an integrated hardware and software stack, robust and accurate visual SLAM is now possible. However, SLAM is not yet a plug-and-play algorithm as it requires hardware expertise and calibration. For a minimal single camera setup without additional sensing such as an IMU, in-the-wild SLAM that provides both accurate poses and consistent dense maps does not exist. Achieving such a reliable dense SLAM system would open new research avenues for spatial intelligence.
Performing dense SLAM from only 2D images requires reasoning over time-varying poses and camera models, as well as 3D scene geometry. To solve such an inverse problem of large dimensionality, a variety of priors, from handcrafted to data-driven, have been proposed. Single-view priors, such as monocular depth and normals, attempt to predict geometry from a single image, but these contain ambiguities and lack consistency across views. While multi-view priors like optical flow reduce ambiguity, decoupling pose and geometry is challenging since pixel motion depends on both the extrinsics and the camera model. Although these underlying causes may vary across time and different observers, the 3D scene remains invariant across views. Therefore, the unifying prior required to solve for poses, camera models, and dense geometry from images is over the space of 3D geometry in a common coordinate frame.
Figure 1: Reconstruction from our dense monocular SLAM system on the Burghers sequence [56]. Using two-view predictions from MASt3R shown on the left, our system achieves globally consistent poses and geometry in real-time without a known camera model.
@@ 2 Related Work
All systems mentioned thus far assume known intrinsic calibration. Classical automatic intrinsic calibration is possible when there are strict assumptions on scene geometry or unchanging parameters across a set of images [13], but encounters degenerate configurations and sensitivity to noise. Given an initial estimate of intrinsics, refinement via bundle adjustment can improve accuracy online [17], but this already assumes a parametric model and sufficient initialisation of all parameters. Combining DROID-SLAM and self-calibration [12] yields improved robustness to noisy intrinsics, but optimisation is slower due to denser matrix fill-in. More recently, data-driven methods predict intrinsics from one or multiple images [14, 48], but are either limited in accuracy for in-the-wild SLAM or are not flexible in the camera model definition.
@@ 3.2 Pointmap Matching
By using the nonlinear least-squares form similar to [35], we can iteratively solve for updates to projected locations by calculating analytical Jacobians and solving via Levenberg-Marquardt. This can be done separately for each point and converges for almost all valid pixels within 10 iterations as the ray image is smooth. At the end of this process, we now have initial matches ${{\mathbf{m}}}_{i,j}$. When there is no initial estimate for the projection $\mathbf{p}$, such as when tracking against a new keyframe or when matching loop closure edges, all pixels are initialised with the identity mapping. During tracking, since we always have the matches from the previous frame, we can use this as initialisation to further speed up the convergence. To handle occlusions and outliers, we also invalidate matches that have large distances in 3D space. Our matching is massively parallel on GPU and additionally can leverage the incremental nature of SLAM.
@@ 3.3 Tracking and Pointmap Fusion
While ${{\mathbf{X}}}^{k}_{f}$ instead of ${{\mathbf{X}}}^{f}_{f}$ could also be aligned to ${{\mathbf{X}}}^{k}_{k}$ with the benefit of no explicit matching required as they are pixel aligned, we found that explicit matching with ${{\mathbf{X}}}^{f}_{f}$ had improved accuracy for larger baseline scenarios. More importantly, although the 3D point error is suitable, it is easily skewed by errors in the pointmap predictions as inconsistent predictions in depth are relatively frequent. Since we ultimately fuse predictions into a single pointmap that averages out all the predictions, error in tracking degrades the keyframe’s pointmap that will also be used in the backend.
@@
This results in a similar angular error as mentioned in Eq. 3 and shown in Fig. 2, except that we now have many known correspondences and wish to find the pose that minimises all angular errors between canonical rays and corresponding predicted rays from the current frame. Since angular errors are bounded, ray-based errors are robust against outliers [30]. We also include an error term with a small weight on the difference in distances from the camera centre. This prevents the system from becoming degenerate under pure rotation, while avoiding significant bias from errors in depth. We efficiently solve for updates to the pose using Gauss-Newton in an iteratively reweighted least-squares (IRLS) framework. We calculate analytical Jacobians of the ray and distance errors with respect to a perturbation ${\boldsymbol{\tau}}$ of the relative pose ${{\mathbf{T}}}_{kf}$. We stack the residuals, Jacobians, and weights into matrices $\mathbf{r}$, $\mathbf{J}$, and $\mathbf{W}$, respectively. We iteratively solve the linear system and update the pose via:
@@
The pointmap initially has larger errors and less confidence due to only using small baseline frames, but filtering merges information from many viewpoints. We experimented with different ways of updating the canonical pointmap, and found that weighted average was best for maintaining coherence while filtering out noise. Compared to the canonical pointmap in MASt3R-SfM [10], we compute this incrementally and require transformation of the points since an additional network prediction of ${{\mathbf{X}}}^{k}_{k}$ would slow down tracking. Filtering has a rich history in SLAM, and yields the benefit of leveraging information from all frames without having to explicitly optimise for all camera poses and store all predicted pointmaps from the decoder in the backend.
@@ 3.4 Graph Construction and Loop Closure
When tracking, a new keyframe $\mathcal{K}_{i}$ is added if the number of valid matches or the number of unique keyframe pixels in ${{\mathbf{m}}}_{f,k}$ falls below a threshold $\omega_{k}$. After adding $\mathcal{K}_{i}$, a bidirectional edge to the previous keyframe $\mathcal{K}_{i-1}$is added to the edge-list $\mathcal{E}$. This constrains the estimated poses sequentially in time; however, drift can still occur. To close both small and large loops, we adapt the Aggregated Selective Match Kernel (ASMK) [46, 47] framework used by MASt3R-SfM [10] for image retrieval from encoded features. While this was previously used in a batch setting where all images are available from the start, we modify it to work incrementally. We query the database with the encoded features of $\mathcal{K}_{i}$ to obtain the top-$K$ images. Since the codebook only has tens of thousands of centroids, we found that a dense L2 distance calculation was sufficiently fast to quantise the features. If the retrieval scores are above a threshold $\omega_{r}$, we give these pairs to the MASt3R decoder and add bidirectional edges if the number of matches from Sec. 3.2 is above a threshold $\omega_{l}$. Lastly, we update the retrieval database by adding the new keyframe’s encoded features to the inverted file index.
@@ 3.5 Backend Optimisation
where ${{\mathbf{T}}}_{ij}={{\mathbf{T}}}_{WC_{i}}^{-1}{{\mathbf{T}}}_{WC_{j}}$. Given N keyframes, Eq. 9 forms and accumulates $14\times 14$ blocks into the $7N\times 7N$ Hessian. We solve this problem again using Gauss-Newton as in Eq. 7 but with sparse Cholesky decomposition as the system is not dense. Construction of the Hessian is made efficient through the use of analytical Jacobians and parallel reductions all implemented in CUDA. Again, a small error term on consistency in distances is added to avoid degeneracy in the pure-rotation case. At most 10 iterations of Gauss-Newton are performed for every new keyframe and optimisation terminates upon convergence. Second-order information greatly speeds up the global optimisation over the alternatives, and our efficient implementation ensures that it is not the bottleneck in the overall system.
@@ 3.6 Relocalisation
If the system loses tracking due to an insufficient number of matches, relocalisation is triggered. For a new frame, the retrieval database is queried with a stricter threshold on the score. Once the retrieved images have a sufficient number of matches with the current frame, it is then added as a new keyframe into the graph and tracking resumes.
@@ 4.1 Camera Pose Estimation
TUM RGB-D: On the TUM dataset, we demonstrate state-of-the-art trajectory error when using calibration as shown in Tab. 1. Many of the previously best performing algorithms, such as DROID-SLAM, DPV-SLAM, and GO-SLAM, build on the foundational matching and end-to-end system proposed by DROID-SLAM. In contrast, we propose a unique system that takes an off-the-shelf two-view geometric prior and show that it can outperform other systems while operating in real-time. Furthermore, our uncalibrated system significantly outperforms a baseline, which we denote DROID-SLAM*, that calibrates the intrinsics using GeoCalib [48] on the first image of a sequence, which is then used by DROID-SLAM. We achieve this without assuming a fixed camera model across the entire sequence, and demonstrate the value of 3D priors for dense uncalibrated SLAM over priors that solve subproblems. Our uncalibrated SLAM results are also comparable to results from recent learned techniques such as DPV-SLAM with known calibration.
@@ 4.2 Dense Geometry Evaluation
7-scenes ATE Accuracy Completion Chamfer
DROID-SLAM 0.049 0.115 0.040 0.077
Spann3R @20 N/A 0.069 0.047 0.058
Spann3R @2 N/A 0.124 0.043 0.084
Ours 0.047 0.074 0.057 0.066
Ours* 0.066 0.068 0.045 0.056
EuRoC ATE Accuracy Completion Chamfer
DROID-SLAM 0.022 0.173 0.061 0.117
Ours 0.041 0.099 0.071 0.085
Ours* 0.164 0.108 0.072 0.090
@@ 4.3 Qualitative Results
4.4 Component Analysis
We compare matching techniques in Tab. 5. Our parallelised projective matching with feature refinement achieves the best accuracy with significantly faster runtime. Performing MASt3R matching over all pixels takes 2 seconds, while our matching takes 2ms and makes the entire system FPS nearly 40x faster. Please refer to the supplementary for a full runtime analysis of the system. In Tab. 5, we test different methods for updating the canonical pointmap and report the average ATE across TUM, 7-Scenes, and EuRoC. Selecting the most recent and first pointmaps incur drift and lack sufficient baseline, respectively. Given calibration, weighted fusion performs on par with selecting the pointmap with the highest median confidence, but it achieves the lowest ATE without calibration and improves the ATE on EuRoC by 1.3cm, indicating that fusing over camera models is important. In Tab. 6, the ray error formulation for uncalibrated tracking and backend optimisation improves performance over using the 3D point error which contains inaccurate depth predictions. Tab. 7 shows that loop closure improves both pose and geometry accuracy, with more significant gains on longer sequences. This demonstrates that the outputs of MASt3R still contain bias and cause drift, which our components are designed to mitigate.
Table 4: Matching comparison.
ATE [m] ATE [m] Matching System
w/ calib w/o calib Time [ms] FPS
$k$-d tree 0.061 0.115 40 8.8
MASt3R 0.042 0.098 2000 0.4
Ours 0.062 0.092 0.5 15.1
Ours + feat 0.039 0.097 2 14.9
Table 5: Fusion methods.
ATE [m] ATE [m]
@@ 4.4 Component Analysis
Median 0.102 0.039
Weighted 0.097 0.039
Table 6: ATE (m) for error formulation in format (point / ray).
TUM 7-Scenes EuRoC avg
0.092 / 0.060 0.084 / 0.066 0.290 / 0.164 0.155 / 0.097
Table 7: Loop closure ablation in format (without LC / with LC).
ATE (m) Chamfer (m)
TUM 7-scenes EuRoC Vicon EuRoC Vicon
Calib 0.064 / 0.030 0.066 / 0.047 0.233 / 0.029 0.151 / 0.085
No Calib 0.090 / 0.060 0.075 / 0.066 0.349 / 0.122 0.179 / 0.090
@@ 6 Conclusion
7 Acknowledgement
This research is supported by the Engineering and Physical Sciences Research Council [grant number EP/W524323/1].
@@ 12 EuRoC Results
13 Comparison to Other SLAM/SfM Methods
13.1 DROID and DPV SLAM
Our system uses a two-view geometric prior in a modular system, while DROID and DPV SLAM learn a matching prior as part of an end-to-end system with differentiable bundle adjustment. While these systems are very accurate for pose estimation, there are fundamental limitations for geometry and generality.
First, bundle adjustment cannot guarantee coherent geometry even with accurate poses, as it lacks smoothness regularisation and constraints under low parallax. Given MASt3R, we find local fusion and scale optimisation to be sufficient for consistency and coherence, while the BA of DROID loses the latter. Second, beyond improved geometry, geometric priors enable new capabilities like continuously changing intrinsics. DROID fixes the model to pinhole during training, and also cannot efficiently handle time-varying intrinsics as this slows down backend optimisation.
13.2 MASt3R-SfM
MASt3R-SfM uses sparse correspondences (subsampling 1/64 pixels) due to MASt3R’s brute-force matching. Our dense projective pointmap matching formulates search as local optimisation and achieves a 1000x speedup without compromising accuracy as shown in Tab. 5. Global optimisation in MASt3R-SfM uses a 1st-order optimiser, lacks minimal rotation updates, and introduces degenerate solutions that require scale renormalisation. To avoid such problems, we formulate a nonlinear least-squares problem and develop a 2nd-order optimiser with minimal pose updates and gauge fixing. Our uncalibrated ray formulation achieves a similar accuracy as MASt3R-SfM’s procedure of fitting pinhole models and minimising reprojection error, but we avoid selecting a specific camera model. This maintains generality of our SLAM system in order to handle all types of distortion, such as fisheye in the future.